OpenStudy (anonymous):

what is the derivitive of e^(x^5-4x)

OpenStudy (anonymous):

\[e ^{x^5-4x}(5x^4-4)\]

OpenStudy (anonymous):

what would the second derivitive be?

OpenStudy (anonymous):

use the product rule and chain rule on branlegr's equation

OpenStudy (anonymous):

hmmm tried that... must not be typing right

OpenStudy (anonymous):

e^(x^5-4x)*5x^4-4)

OpenStudy (anonymous):

\[e^{x^5-4x}(20x+(5x^4-4)^2)\]

OpenStudy (anonymous):

you dont need the product rule, all you need is chain

OpenStudy (anonymous):

derivative of the outside times the derivative of the inside

OpenStudy (anonymous):

e^x's derivative is e^x so consider x = x^5-4x

OpenStudy (anonymous):

which becomes your inside function that you take the derivative and multiply it by

OpenStudy (anonymous):

\[e^{x^5−4x}(20x^3+(5x^4−4)^2)\] I made a mistake... this is it

OpenStudy (anonymous):

thats wrong. You are making it more complicated then it has to be Its just the derivative times the inside. Hes just entering it wrong Im pretty sure

OpenStudy (anonymous):

nadeen is right ... thank you much

OpenStudy (anonymous):

derivative of the outside times the derivative of the inside i mean*

OpenStudy (anonymous):

No, you must use the product rule to find second derivative.

OpenStudy (anonymous):

why are you taking the second derivative, Is that the problem?

OpenStudy (anonymous):

yea nadeem is right, I didnt realize you needed the second derivative

OpenStudy (anonymous):

Thanks....... for a second I thought I was going crazy

OpenStudy (anonymous):

yes higher order

OpenStudy (anonymous):

you never specified the problem correctly creeg no Nadeem, you are 100% correct

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

ln(15x7)−3y3=x3ey... guys want to help with thsi one .. i just need y prime.. cant get a right answer... tried like 1 hr on this stupid problem..

OpenStudy (anonymous):

15x^7

OpenStudy (anonymous):

I thought i did that already

OpenStudy (anonymous):

no this is just an implicit problem

OpenStudy (anonymous):

ln(x)= 1/x is the derivative umm -9y^2 y' is that derivative 3x^2 ey wont you have to do product rule

OpenStudy (anonymous):

sorry this is it.. ln(15x^7)-3y^3=(x^3)(e^y)

OpenStudy (anonymous):

\[(7−3x^3e^y)/x(9y^2+x^3e^y)=dy/dx\] is this not it?

OpenStudy (anonymous):

just isolate y' this is easy

OpenStudy (anonymous):

yes i know what to do saif... im just having a hard time either typing it in or im missing something

OpenStudy (anonymous):

is this webassign?

OpenStudy (anonymous):

cause webassign always messes up everything. I feel your pain

OpenStudy (anonymous):

webwork .. prob the same.. i hate it

OpenStudy (anonymous):

yea i bet!

OpenStudy (anonymous):

and nadeen thats not the answer

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

I dont understand Saif623 check my answer out and see if it makes sence...... you are suppose to differentiate implicitly... right?

OpenStudy (anonymous):

my bad nadeen .. your right... i had to just fix a typo..

OpenStudy (anonymous):

Man..........yall are trippin!

OpenStudy (anonymous):

yea i got the same nadeem

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

lol... its been a long week.. how about tangent lines? e^(yx^3)−x(y^3+1)=9 at (-8,0)

OpenStudy (anonymous):

i have finals on friday and I am doing your hw

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

same here Bro...wtf... lol

OpenStudy (anonymous):

that one will make my head explode with all my other studying im doing right now

OpenStudy (anonymous):

sorry man you are on your own

OpenStudy (anonymous):

ya i just got it .. nevermind.. thanks a lot you two.. big help... good luck with the finals..

OpenStudy (anonymous):

no problem and Thanks Im gonna need it. 3 finals back to back

OpenStudy (anonymous):

ya thats my schedule tomarrow..

OpenStudy (anonymous):

Professors are so annoying. I have a webassign due to. Thank GOD I asked for an extension so its due after my finals

OpenStudy (anonymous):

my math prof gave a quiz yesterday and this hw is do right now and the test is tomarrow night.. bad planning on their part .. whole class is pissed..

OpenStudy (anonymous):

i bet! my professor randomly gives webassign homework so we will have like an hour to do it. Gets annoying

OpenStudy (anonymous):

Lol when I see people asking on here problems to do with adding fractions. I wish that was my hw lol!

OpenStudy (anonymous):

damn that would be frustrating.. lol ya ... alright well thanks again.. i got some more studying to do so ill cya around on here .. late

OpenStudy (anonymous):

lates I got a lot of studying ahead of me to. Good Luck with the exam tomorrow

OpenStudy (anonymous):

thanks ...